# {centering}Tests of restricted Quantum Focusing and a universal CFT bound

### v1

@@ Abstract
Second, for $d>2$, we derive a new CFT constraint: the coincident limit of the product of two averaged null energy operators in any CFT cannot be zero. Equivalently, the rQFC imposes an upper bound of 2d12d-1 on the scaling dimension of the leading even-spin Regge trajectory, analytically continued in $J$ and evaluated at $J=3$. The bound holds in free and weakly-interacting CFTs, and is saturated by the double-trace trajectories in planar $\mathcal{N}=4$ Super Yang-Mills at strong coupling. We speculate about a strengthened Quantum Null Energy Condition that would imply our results.
@@ 1 Introduction
In Sec. 4, we show that the rQFC implies new constraints purely in CFT: a novel upper bound on the scaling dimension of the leading operators contributing to the Operator Product Expansion (OPE) of two averaged null energy operators. This is a new instance of quantum gravity constraints predicting QFT constraints. A well-known previous example is the QNEC, which was originally derived in [1] from the QFC, and subsequently proved in several papers [17, 18, 19, 20, 21].
@@
In Sec. 4, we find that a different $G\to 0$ limit of the rQFC, one in which we simultaneously send $\Sigma\to 0$ while keeping $\Sigma^{d-2}\gg G$ (as required for validity of perturbation theory) gives a new (stronger) bound on CFTs. [^13] For instance, let $\Sigma=L^{1-\alpha}G^{\frac{\alpha}{d-2}}$ for $0<\alpha<1$, where $L$ is some large scale associated to the state under consideration. This is because the limit creates a competition between the first and the second lines on the RHS of $\partial_{\lambda}\Theta_{\lambda}$ in Eq. (1.14). We work this out explicitly in a class of near-vacuum states where we have a lot of computational control over the terms in Eq. (1.14), and extract from the rQFC a certain upper bound on the dimension of the leading even-spin Regge trajectory in CFTs. We sketch the basic idea and main results in the rest of this section, and we also speculate about a more general stronger-than-QNEC bound.
Figure 2: Application of the rQFC to the domain of dependence $\mathcal{W}$ of a ball in Minkowski space (left). We consider null deformations of the edge $\eth\mathcal{W}$, parametrized by $y^{a}$, to $v=V_{\lambda}(y^{a})$ with width $\Sigma$. A new non-gravitational limit with $G,\Sigma\to 0$ and $\Sigma^{d-2}\gg G$ yields a CFT bound. As part of the derivation, $\mathcal{W}$ is conformally mapped to a Rindler wedge $\tilde{\mathcal{W}}$ with deformation $\tilde{v}=\tilde{V}_{\lambda}(\tilde{y}^{a})$, where von Neumann entropy derivatives relate to the coincident limit of two $\mathcal{E}$ insertions (red), governed by their OPE (right). The bound constrains the leading operator in the $\mathcal{E}\times\mathcal{E}$ OPE.
@@
For a generic class of near-vacuum states (i.e., close to one inner product with the vacuum), it turns out that the Rindler wedge quantity on the RHS of Eq. (1.17) is suppressed by $\Sigma^{\delta_{0}-d-1}$, where the theory-dependent number $\delta_{0}$ will be described momentarily [37] (see also [38], and Appendix B for a review). Using this result in Eq. (1.14), we obtain in a CFT:
@@
$\begin{split}\Theta_{\lambda}(y^{a})=0\implies\partial_{\lambda}\Theta_{\lambda}(y^{a})=\frac{16(d-1)}{d-2}G^{2}\left(\frac{1}{\sqrt{h_{\lambda}}}\left.\frac{\delta S_{\rm ren}}{\delta V(y^{a})}\right\rvert_{\lambda}\right)^{2}+O(G\Sigma^{\delta_{0}-d-1})+\cdots.\end{split}$ (1.18)
In Sec. 4 and Appendix C, we argue that the omitted terms (in the ellipsis) are $o(G^{2})$ and therefore unimportant for the analysis.
In Eq. (1.18), $\delta_{0}$ is the scaling dimension of the leading operator contributing to the so-called light-ray OPE of the averaged null energy operators $\mathcal{E}$:
$\mathcal{E}(\tilde{y}^{a})=\int_{-\infty}^{\infty}d\tilde{v}~T_{ij}(\tilde{v},\tilde{y}^{a})\tilde{k}^{i}\tilde{k}^{j},$ (1.19)
where $\tilde{v}$ is an affine parameter of an infinite null line, $\tilde{k}^{i}=(\partial_{\tilde{v}})^{i}$, and $\tilde{y}^{a}$ specify the transverse location of the geodesic on the Rindler horizon.
The $\mathcal{E}\times\mathcal{E}$ OPE can be expressed as follows:
$\mathcal{E}(0)\mathcal{E}(\tilde{y}^{a})\stackrel{{\scriptstyle|\tilde{y}|\to 0}}{{=}}$ $\sum_{j,n}c_{j,n}~|\tilde{y}|^{\delta_{j,n}-2(d-1)}\hat{y}^{a_{1}}\cdots\hat{y}^{a_{j}}~\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}(0)+\text{descendents of $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$}$, (1.20)
where $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$, with conformal dimension $\delta_{j,n}$, are the so-called light-ray operators with transverse spin $j$, and associated to the $n$-th even-spin Regge trajectory (see Fig. 10). [^14] Light-ray operators are non-local and can be defined, roughly speaking, through smearing of two local CFT operators in a certain way in a neighborhood of a null line. A thorough discussion of these operators and their OPE structure is beyond the scope of this paper, but can be found in [39, 40, 41, 42] In particular, the spectrum $\delta_{j,n}$ can be phrased in terms of that of local CFT operators as follows:
$\delta_{j,n}=$ $-1+\Delta^{\text{even}}_{j,n}(J=3),~~~j=0,2,4.$ (1.21)
$\delta_{j,n}=$ $-1+\Delta^{\text{even}}_{4,n}(J=3+2p),~~~j=4+2p,~(p\in\mathbb{Z}^{+}),$ (1.22)
where $\Delta^{\text{even}}_{j,n}(J)$ is the scaling dimension of the $n$-th even-spin Regge trajectory associated to the transverse (along $\eth\mathcal{W}$) spin $j$ sector, analytically continued in total spin JJ. [^15] Note that $J$ and $j$ label the symmetric traceless representation of SO(d)(d) and SO(d2)(d-2) respectively.
The $\Sigma$ scaling of the RHS of Eq. (1.17) (as $\Sigma\to 0$) is governed by the leading term (as $|\tilde{y}|\to 0$) in the OPE (1.20) evaluated in a certain matrix element associated to our near-vacuum state [37]. The leading term in this OPE is then controlled by the leading ($n=1$) Regge trajectory. Let us define:
$\delta_{0}\equiv\text{min}_{j}\delta_{j,n=1}.$ (1.23)
From Eq. (1.14), we can deduce that the rQFC would be violated if $\delta_{0}$ is too large, as that would suppress the second term on the RHS of $\partial_{\lambda}\Theta_{\lambda}$ in Eq. (1.18) compared to the positive first term. This argument requires a simultaneous $G,\Sigma\to 0$ limit which, for the sake of remaining within our perturbative regime, must satisfy
@@
It is easy to see that if the power of $\Sigma$ in the second term on the RHS of Eq. (1.18) is larger than d2d-2, it is possible to make $\partial_{\lambda}\Theta_{\lambda}$ negative while staying within the regime (1.24). Therefore, the rQFC implies that in any CFT:
$\delta_{0}\leq 2(d-1).$ (1.25)
From the OPE (1.20), we see that an equivalent way to state this bound is:
$\lim_{|\tilde{y}|\to 0}\mathcal{E}(0)\mathcal{E}(\tilde{y}^{a})\neq 0.$ (1.26)
Since we are free to choose our near vacuum state with an operator of our choice, if our theory contains superselection sectors, the statements (1.25) and (1.26) are implied within each sector.
To our knowledge, this is a novel bound on the spectrum of CFT operators. Note in particular that using Eqs. (1.21) and (1.22), the bound directly constrains the spectrum of local operators, again within each superselection sector. We can summarize our results in the following diagram:
It is easy to see that our bound is obeyed by free theories, since there the leading Regge trajectory satisfies $\text{min}_{j}\Delta^{\text{even}}_{j,1}(J)=d-2+J$, and therefore through Eq.(1.21) we learn that $\delta_{j,1}=d$, satisfying our bound (1.25) with room to spare. Therefore, we expect that weakly coupled CFTs would also satisfy the bound. For a strongly coupled example, it is interesting to study the case of $\mathcal{N}=4$ SU(N)(N) Super Yang-Mills. [^16] We thank David Simmons-Duffin for explaining to us the story of $\mathcal{N}=4$ Super Yang-Mills. In the $N\to\infty$, limit, we can delineate the Regge trajectories of single and double-trace operators. It turns out that in the large t’Hooft coupling limit, the single-trace $\text{min}_{j}\Delta^{\text{even}}_{j,1}(J)$ can be arbitrarily increased for each $J$, but the leading double-trace trajectory satisfies $\text{min}_{j}\Delta^{\text{even}}_{j,1}(J)=4+J$ (approaching it from below in the $N\to\infty$ limit) [41]. Therefore, the double-trace trajectory leads to $\text{min}_{j}\delta_{j,1}=6$, which saturates the bound (1.25) given d=4d=4.171717For instance, one can consider the trajectory associated with the stress-tensor double-twist $TT]_{J}$, which at large spin have twist (i.e., scaling dimension minus spin) $\tau(J)=2d+J-(J+4)=2d-4$. We then obtain $\Delta(J=3)=\tau(J=3)+3=2d-1$.
Since the constraint (1.25) is a pure CFT statement, one can attempt to prove it directly, similarly to how the QNEC was proved after being predicted by gravity. We leave this to future work.
Let us end the introduction with a speculation. Our new CFT bound is stronger than the QNEC in the particular class of near-vacuum states we consider in Sec. 4. But it is unsatisfactory to only have such a bound in this limited regime. It is natural to expect that, in fact, a universal (holding in all QFTs in all states) stronger-than-QNEC bound is true. In the $\Sigma\to 0$ regime, this bound would presumably prevent the (positive) LHS of Eq. (1.15) from scaling with $\Sigma^{\alpha}$ with $\alpha>d-2$, whenever $\delta S_{\rm ren}/\delta V(y^{a})\neq 0$. These conditions can culminate in the following natural bound for null deformations of the Rindler wedge:
@@
where $\kappa>0$ is to be determined. This bound would then hold by either the LHS scaling with $\Sigma^{\alpha}$ as $\Sigma\to 0$ for $0\leq\alpha<d-2$, or with $\alpha=d-2$ in which case the competition of magnitudes between LHS and RHS becomes non-trivial. We leave an exploration of this bound to future work.
@@ 3 The rQFC and QFC in JT gravity coupled to a QFT
In this section, we test the rQFC in a particular setting in $d>2$ perturbative semiclassical gravity regime. We discover that the rQFC implies a novel constraint in all CFTs: [^29] The reader may wonder why instead of all CFTs, we do not demand the new constraint only of CFTs which arise as low energy limits of consistent UV-complete theories of quantum gravity. Historically, similar gravitational constraints, like the generalized second law (GSL), end up implying things about the low energy QFT sector (monotonicity of relative entropy in the case of the GSL) which are true in general QFT [44]. So demanding the validity of the new constraint in all CFTs is informed by this precedent. the OPE of two averaged null energy operators in Minkowski spacetime is either singular or approaches a non-zero constant. Equivalently, the constraint can also be stated as an upper bound on the (analytically continued in spin) dimensions of CFT Regge trajectories. To obtain this result, we make crucial use of [37] which relates von Neumann entropy of null deformed Rindler wedges to light-ray operators. We end by some speculative stronger bounds, including a universal strengthened QNEC.
@@
In [37], the LHS of Eq. (4.24) was calculated in d>2d>2.343434In fact, the expression which we cite is true for any QFT, not just CFTs. We describe the setup, and present this expression from [37], and relegate a sketch of its derivation to Appendix B.
@@
$\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}^{(1)}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda}=-\frac{\epsilon^{2}}{2}\int d^{d-2}\tilde{y}^{\prime}\sqrt{\tilde{h}}~\partial_{\lambda}\tilde{V}_{\lambda}(\tilde{y}^{\prime a})~\int_{-\infty}^{\infty}ds~e^{s}\langle\tilde{O}^{(1)}_{f}\mathcal{E}(\tilde{y}^{\prime a})\mathcal{E}(\tilde{y})\tilde{O}^{(1)}_{f_{s}}\rangle+O(\epsilon^{3}),$ (4.27)
@@
where $T_{\tilde{v}\tilde{v}}$ denotes the indicated component of the stress-energy tensor, and $\tilde{O}^{(1)}_{f_{s}}$ is a result of boosting $\tilde{O}^{(1)}_{f}$ by an amount proportional to ss. [^35] More accurately, $O_{f_{s}}$ must denote the modular-evolved by $s$ version of operator $O_{f}$. But if we evaluate our expression at $\lambda=0$, the corresponding modular operator is simply a boost. For the rest of our argument, we can assume everything is being evaluated at $\lambda$, or that $O_{f_{s}}$ is modular-evolved in general (see Appendix B). The distinction is irrelevant for our results. This can be arranged by a change of the smearing function $f$ to $f_{s}$:
$f_{s}(x)=f(xe^{s}).$ (4.29)
We are interested in deriving how the LHS of Eq. (4.27) scales with the parameters involved. Specifically, let us discuss how the RHS term at $O(\epsilon^{2})$ scales with parameter the transverse width of the deformation $\Sigma$ and the scale $L$ associated to the state (as specified by the size of the smearing function in Eq. (4.27)). The dependence on $\Sigma$ comes through the transverse integral on the RHS of Eq. (4.27). In particular, the behavior of the integrand in Eq. (4.27) as $\tilde{y}$ approaches zero determines the magnitude of the LHS at $O(\epsilon^{2})$. The $\tilde{y}\to 0$ structure of the integrand is in turn controlled by the operator product expansion (OPE) of averaged null energy operators. This is a less conventional OPE which involves light-ray operators (i.e., operators with support on null generators of the $\tilde{u}=0$ plane) as they approach each other in the transverse direction. The averaged null energy operator $\mathcal{E}$ is an example of a light-ray operator. See [41, 42] for a proper discussion of light-ray operators and their OPEs.
Figure 10: The OPE of two averaged null energy operators $\mathcal{E}(\tilde{y}^{a})=\int_{-\infty}^{\infty}d\lambda~T_{\tilde{v}\tilde{v}}(\tilde{v}=\lambda,\tilde{y}^{a})$ is shown. The RHS of the OPE consists of so-called light-ray operators $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$, labeled by a transverse spin $j$, and corresponding to a certain $n$-th (even spin) Regge trajectory. In our states, the second derivative of the von Neumann entropy under the null deformations of the Rindler wedge are controlled by this OPE: the more singular (as $|\tilde{y}|\to 0$) the OPE, the smaller the second derivative becomes in the limit of deformation width going to zero. The rQFC enforces an upper bound on the scaling of this second derivative with the null-deformation width, which can be translated into a the upper bound (4.38) on the dimensions $\delta_{j,n}$.
@@
$\mathcal{E}(0)\mathcal{E}(\tilde{y}^{a})\stackrel{{\scriptstyle|\tilde{y}|\to 0}}{{=}}$ $\sum_{j,n}c_{j,n}~|\tilde{y}|^{\delta_{j,n}-2(d-1)}\hat{y}^{a_{1}}\cdots\hat{y}^{a_{j}}~\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}(0)+\text{descendents of $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$}$, (4.30)
where $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$ denotes a so-called light-ray operator with scaling dimension $\delta_{j,n}$, and transverse spin $j$ (see Fig. 10). Each operator $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$ is a non-local operator, and could be defined through smearing a pair local CFT operators in a certain way along a null geodesic. We will not cover the details of the light-ray OPE story here, but its structure (4.30) is sufficient for us to extract a certain bound on the spectrum $\delta_{j,n}$. Conveniently, the spectrum $\delta_{j,n}$ could be described simply in terms of the spectrum of local CFT data, allowing us the ability to phrase the bound in terms of data of the local CFT spectrum. Let $J$ denote total spin (as opposed to transverse spin $j$), and let $\Delta^{\text{even}}_{j,n}(J)$ be the analytically continued in $J$ of the scaling dimension of the n-th Regge trajectory in the transverse spin $j$ sector. We then have [40, 41, 42]:
$\delta_{j,n}=$ $\Delta^{\text{even}}_{j,n}(J=3)-1,~~~j=0,2,4.$ (4.31)
$\delta_{j,n}=$ $\Delta^{\text{even}}_{j=4,n}(J=3+2p)-1,~~~j=4+2p,~(p\in\mathbb{Z}^{+}).$ (4.32)
It is easy to see from the OPE (4.30), that the leading $\Sigma\to 0$ behavior of Eq. (4.27) is given by the smallest $\delta_{j,n}$. Let
$\delta_{0}\equiv\text{min}_{n,j}\delta_{j,n}.$ (4.33)
In particular, in the limit of small $\Sigma$ we have:
$-\epsilon^{2}\int_{0}^{\Sigma}d\tilde{\rho}~\tilde{\rho}^{d-3}$ $\exp\left[\frac{1}{1-\Sigma^{2}/\tilde{\rho}^{2}}\right]\tilde{\rho}^{\delta_{0}-(2d-1)}$
$\times\int d\tilde{V}_{d-3}ds~e^{s}~\hat{y}^{a_{1}}\cdots\hat{y}^{a_{j}}\langle{O}_{f}\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}(0){O}_{f_{s}}\rangle=O\left(\epsilon^{2}\frac{\Sigma^{\delta_{0}-d-1}}{L^{\delta_{0}-1}}\right),$ (4.34)
where $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}(0)$ is the light-ray operator corresponding to $\delta_{0}$, and $dV_{d-3}$ denotes the volume form of a (d3)(d-3)-sphere part of the transverse directions $\eth\tilde{\mathcal{W}}$ (see Eq. (4.19)). Only the $\tilde{\rho}$ integral in Eq. (4) gives rise to non-trivial $\Sigma$ scaling, namely $\Sigma^{\delta_{0}-d-1}$. The $V_{d-3}$ and $s$ integrals will in general be tedious to compute, but by dimensional analysis it must only result in the factors of $L$ in the denominator of Eq. (4). Note that by the QNEC, Eq.(4) is non-positive.
@@
$\frac{4(d-1)}{d-2}\ell_{S}^{2(d-2)}\left(\left.\frac{1}{\sqrt{h_{\lambda}}}\frac{\delta S^{(1)}_{\rm ren}}{\delta V(y^{a})}\right\rvert_{\lambda}\right)^{2}$ $=O\left(\epsilon^{2}\frac{\ell_{S}^{2(d-2)}}{L^{2(d-1)}}\right)>0,$ (4.35)
$\frac{4\ell_{S}^{d-2}}{\sqrt{h_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}^{(1)}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda}$ $=O\left(\epsilon^{2}\frac{\ell_{S}^{d-2}\Sigma^{\delta_{0}-d-1}}{L^{\delta_{0}-1}}\right)<0,$ (4.36)

### v2

# {centering}Tests of restricted quantum focusing and a new CFT bound

@@ Abstract
Second, for $d>2$, we derive from the rQFC a constraint stronger than the Quantum Null Energy Condition (QNEC). In a broad class of states, this bound forbids the QNEC from saturating faster than $O(\mathcal{A})$ as the transverse area $\mathcal{A}$ of a certain null deformation shrinks to zero. We speculate about a universal strengthened QNEC holding across all QFT states.
@@ 1 Introduction
In Sec. 4, we show that in a class of near-vacuum states, the rQFC implies a new CFT bound. This is a new instance of quantum gravity constraints predicting QFT constraints. A well-known previous example is the QNEC, which was originally derived in [1] from the QFC, and subsequently proved in several papers [17, 18, 19, 20, 21]. The new CFT bound is stronger than the QNEC.
@@
In Sec. 4, we find that a different $G\to 0$ limit of the rQFC, one in which we simultaneously send $\Sigma\to 0$ while keeping $\Sigma^{d-2}\gg G$ (as required for validity of perturbation theory) gives a new (stronger) bound on CFTs. [^13] For instance, let $\Sigma=L^{1-\alpha}G^{\frac{\alpha}{d-2}}$ for $0<\alpha<1$, where $L$ is some large scale associated to the state under consideration. This is because the limit creates a competition between the first and the second lines on the RHS of $\partial_{\lambda}\Theta_{\lambda}$ in Eq. (1.14). We work this out explicitly in a class of near-vacuum states where we have a lot of computational control over the terms in Eq. (1.14). We sketch the basic idea and main results in the rest of this section, and we also speculate about a more general stronger-than-QNEC bound.
Figure 2: Application of the rQFC to the domain of dependence $\mathcal{W}$ of a ball in Minkowski space (left). We consider null deformations of the edge $\eth\mathcal{W}$, parametrized by $y^{a}$, to $v=V_{\lambda}(y^{a})$ with width $\Sigma$. A new non-gravitational limit with $G,\Sigma\to 0$ and $\Sigma^{d-2}\gg G$ yields a stronger-than-QNEC bound in CFTs. As part of the derivation, $\mathcal{W}$ is conformally mapped to a Rindler wedge $\tilde{\mathcal{W}}$ with deformation $\tilde{v}=\tilde{V}_{\lambda}(\tilde{y}^{a})$, where von Neumann entropy derivatives relate to the coincident limit of two $\mathcal{E}$ insertions.
@@
We can combine Eq. (1.17) with Eq. (1.14), to obtain:
@@
$\Theta_{\lambda}(y^{a})=0\implies\partial_{\lambda}\Theta_{\lambda}(y^{a})$ $=\frac{16(d-1)}{d-2}G^{2}\left(\frac{1}{\sqrt{h_{\lambda}}}\left.\frac{\delta S_{\rm ren}}{\delta V(y^{a})}\right\rvert_{\lambda}\right)^{2}$
$-G\left(2\pi\langle\tilde{T}_{ij}\rangle\tilde{k}^{i}\tilde{k}^{j}-\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda}\right)+\cdots.$ (1.18)
It turns out that for a generic class of near-vacuum states: [^14] The near-vacuum states are chosen for simplicity of the analysis in Sec. 4 to have no stress-energy tensor in the vicinity of the deformation profile.
[…]
the Rindler wedge quantity on the second line of Eq. (1) can be computed explicitly [37] (see also [38], and Appendix B for a review). In particular, a non-trivial scaling with $\Sigma$ as $\Sigma\to 0$ can be obtained:
$2\pi\langle\tilde{T}_{ij}\rangle\tilde{k}^{i}\tilde{k}^{j}-\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda}=O(\epsilon^{2}\Sigma^{d-2-\delta})$ (1.20)
where $\delta$ depends on both the theory and the state. We will discuss it in more detail below. Furthermore, we expect that generically: [^15] Here, we use that in the $G\to 0$ limit, $\delta S_{\rm ren}/\delta V(y^{a})$ goes to its Rindler wedge counter-part denoted by $\delta\tilde{S}_{\rm ren}/\delta\tilde{V}(\tilde{y}^{a})$.
$\left.\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\frac{\delta\tilde{S}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda}=O(\epsilon).$ (1.21)
which implies that the $h_{\lambda}^{-1/2}\delta S_{\rm ren}/\delta V(y^{a})=O(\epsilon)$.
Let us discuss how Eqs. (1), (1.20), and (1.21) result in a bound stronger than the QNEC. As explained, the regular $G\to 0$ limit of Eq. (1) yields the QNEC. However, Eq. (1.20) in principle allows us to suppress the second line of Eq. (1) with an independent parameter $\Sigma$ compared to the positive first line. Therefore, a simultaneous $G,\Sigma\to 0$ limit must extract a new bound from the rQFC. [^16] In Sec. 4 and Appendix C, we argue that the omitted terms (in the ellipsis) are $o(G^{2})$ and therefore do not enter this analysis. But since $\Sigma$ controls the transverse size of our deformation profile, to stay in the regime of validity of the computation, it is important that we obey:
@@
It is easy to see that if $\delta<0$ in Eq. (1.20), it is possible to make $\partial_{\lambda}\Theta_{\lambda}$ negative while staying within the regime (1.22).
The conclusion is that an appropriate $G,\Sigma\to 0$ limit of the rQFC implies the following stronger-than-QNEC bound in all CFTs:
$\left.\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\frac{\delta\tilde{S}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda}=O(\epsilon)\implies 2\pi\langle\tilde{T}_{ij}\rangle\tilde{k}^{i}\tilde{k}^{j}-\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda}\stackrel{{\scriptstyle\Sigma\to 0}}{{\geq}}O(\epsilon^{2}\Sigma^{d-2})\geq 0.$ (1.23)
where by the notation $\geq O(\epsilon^{2}\Sigma^{d-2})$, we mean that at $O(\epsilon^{2})$, the expression cannot go to zero faster than $\Sigma^{d-2}$ as $\Sigma\to 0$. In particular, this bound is stronger than the QNEC applied to this state, which only demands that the LHS of Eq. (1.20) is positive.
Let us now comment on the origin of the parameter $\delta$ in Eq. (1.20). In [37], it was found that $\delta$ in Eq.(1.20) can be obtained from the following:
[…]
$\text{where,}~~~\mathcal{E}(\tilde{y}^{a})=$ $\int_{-\infty}^{\infty}d\tilde{v}~T_{ij}(\tilde{v},\tilde{y}^{a})\tilde{k}^{i}\tilde{k}^{j}.$ (1.24)
[…]
where $K$ is the boost generator around the Rindler wedge. In Appendix D, we analyze the rQFC implication that $\delta\geq 0$ from the perspective of the operator product expansion of $\mathcal{E}$’s.
We can summarize the story so far in the following diagram:
Let us end the introduction with a speculation. Our new CFT bound (1.23) is stronger than the QNEC in the particular class of near-vacuum states (1.19). But it is unsatisfactory to only have such a bound in this limited regime. It is natural to expect that, in fact, a universal (holding in all QFTs in all states) stronger-than-QNEC bound is true. In the $\Sigma\to 0$ regime, this bound would presumably prevent the QNEC from scaling with $\Sigma^{\alpha}$ with $\alpha>d-2$, whenever $\delta S_{\rm ren}/\delta V(y^{a})\neq 0$. These conditions can culminate in the following natural bound for null deformations of the Rindler wedge:
@@
where $\kappa>0$ is to be determined. This bound would then hold by either the LHS scaling with $\Sigma^{\alpha}$ as $\Sigma\to 0$ for $0\leq\alpha<d-2$, or with $\alpha=d-2$ in which case the competition of magnitudes between LHS and RHS becomes non-trivial. The bound (1.26) is reminiscent of the $d=2$ strengthened QNEC in CFTs, which was proposed in [35], and proved in general states in [34]. We leave an exploration of this bound to future work.
@@ 3 The rQFC and QFC in JT gravity coupled to a QFT
In this section, we test the rQFC in a particular setting in $d>2$ perturbative semiclassical gravity regime. We discover that the rQFC implies a novel stronger-than-QNEC bound in all CFTs. [^28] The reader may wonder why instead of all CFTs, we do not demand the new constraint only of CFTs which arise as low energy limits of consistent UV-complete theories of quantum gravity. Historically, similar gravitational constraints, like the generalized second law (GSL), end up implying things about the low energy QFT sector (monotonicity of relative entropy in the case of the GSL) which are true in general QFT [40]. So demanding the validity of the new constraint in all CFTs is informed by this precedent. To obtain this result, we make crucial use of [37] which relates von Neumann entropy of null deformed Rindler wedges to light-ray operators. We end by speculating about a more universal strengthened QNEC.
@@
In [37], the LHS of Eq. (4.24) was calculated in d>2d>2.333333In fact, the expression which we cite is true for any QFT, not just CFTs. We describe the setup, and present this expression from [37], and relegate a sketch of its derivation to Appendix B.
@@
$\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}^{(1)}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda=0}=-\frac{\epsilon^{2}}{2}\int d^{d-2}\tilde{y}^{\prime}\sqrt{\tilde{h}}~\partial_{\lambda}\tilde{V}_{\lambda}(\tilde{y}^{\prime a})~\int_{-\infty}^{\infty}ds~e^{s}\langle\tilde{O}^{(1)}_{f}\mathcal{E}(\tilde{y}^{\prime a})\mathcal{E}(\tilde{y})\tilde{O}^{(1)}_{f_{s}}\rangle+O(\epsilon^{3}),$ (4.27)
@@
where $T_{\tilde{v}\tilde{v}}$ denotes the indicated component of the stress-energy tensor, and $\tilde{O}^{(1)}_{f_{s}}$ is a result of evolving $\tilde{O}^{(1)}_{f}$ by a global boost operator by an amount $s$:
$\tilde{O}^{(1)}_{f_{s}}=e^{-iKs}\tilde{O}^{(1)}_{f}e^{iKs}.$ (4.29)
where $K$ is the global boost generator. This can be arranged by boosting the profile $f$ to its boosted version $f_{s}$ [^34] For $\lambda\neq 0$, $O^{(1)}_{f_{s}}$ needs to be replaced by $e^{iKs}\tilde{O}^{(1)}_{f}e^{-iKs}$ where $K$ is the full modular Hamiltonian associated to the cut $\lambda$. Our conclusions are easily generalizable to arbitrary $\lambda$ by this replacement.
We are interested in deriving how the LHS of Eq. (4.27) scales with the parameters involved. Specifically, let us discuss how the RHS term at $O(\epsilon^{2})$ scales with $\Sigma$, the transverse width of the deformation profile, and the scale $L$ associated to the state (as specified by the size of the smearing function in Eq. (4.27)). The dependence on $\Sigma$ is controlled by the behavior of the integrand as $\mathcal{E}$’s approach each other. In particular it is easy to verify that:
$\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}^{(1)}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda=0}=O\left(\frac{\epsilon^{2}\Sigma^{d-2-\delta}}{L^{2d-2-\delta}}\right)$ (4.30)
where:
$\int_{-\infty}^{\infty}ds~e^{s}\langle\tilde{O}^{(1)}\mathcal{E}(\tilde{y}^{\prime a})\mathcal{E}(\tilde{y}^{a})\tilde{O}^{(1)}_{f_{s}}\rangle\stackrel{{\scriptstyle\tilde{y}^{\prime}\to\tilde{y}}}{{=}}O(|\tilde{y}-\tilde{y}^{\prime}|^{{-\delta}})$ (4.31)
@@
$\frac{4(d-1)}{d-2}\ell_{S}^{2(d-2)}\left(\left.\frac{1}{\sqrt{h_{\lambda}}}\frac{\delta S^{(1)}_{\rm ren}}{\delta V(y^{a})}\right\rvert_{\lambda=0}\right)^{2}$ $=O\left(\epsilon^{2}\frac{\ell_{S}^{2(d-2)}}{L^{2(d-1)}}\right)>0,$ (4.32)
$\frac{4\ell_{S}^{d-2}}{\sqrt{h_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}^{(1)}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda=0}$ $=O\left(\epsilon^{2}\frac{\ell_{S}^{d-2}\Sigma^{d-2-\delta}}{L^{2d-2-\delta}}\right)<0,$ (4.33)
where in the first line the generic condition:
$\left.\frac{1}{\sqrt{h_{\lambda}}}\frac{\delta S^{(1)}_{\rm ren}}{\delta V(y^{a})}\right\rvert_{\lambda=0}=\left.\frac{1}{\sqrt{h_{\lambda}}}\frac{\delta\tilde{S}^{(1)}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda=0}=O\left(\frac{\epsilon}{L^{d-1}}\right)$ (4.34)
was used. The first equality in Eq. (4.34) simply comes from the conformal transformation. And the second the inequality is the result of the QNEC, which in this particular setting, follows directly from Eq. (4.27).